OpenStudy (anonymous):

solve y=x^3-9x determine the intercepts? how do I solve for x?

OpenStudy (anonymous):

you mean solve \[x^3-9x=0\] for \(x\) ?

OpenStudy (anonymous):

yes i have to find the intercepts

OpenStudy (zubhanwc3):

well. other than 0 there are no real roots.

OpenStudy (anonymous):

lol factor as \[x(x^2-9)=0\] then as \[x(x+3)(x-3)=0\] and it should be easy from there

OpenStudy (anonymous):

ohh ok thanks!

OpenStudy (zubhanwc3):

x(x^2+9) x= 0 X^2+9 = 0 x^2 = -9 x= +-3i

OpenStudy (zubhanwc3):

plz give me a medal and close ^^

OpenStudy (anonymous):

@zubhanwc3 you misread the question

OpenStudy (zubhanwc3):

@satellite73 no i didn't. solving for x means the x intercepts, i told him, 0, 3i and -3i.

OpenStudy (anonymous):

take a look at the equation

OpenStudy (zubhanwc3):

o.O mb

OpenStudy (zubhanwc3):

your rite:P